# Hyperion: Hierarchical Scheduling for Parallel LLM Acceleration in Multi-tier Networks

### v1

@@ Abstract
LLMs are increasingly executed in edge where limited GPU memory, and heterogeneous compute jointly constrain deployment and motivate model partitioning and request scheduling. In this setting, achieving low latency depends on both model placement and the scheduling of incoming requests across heterogeneous nodes. These two problems are tightly coupled, as a suboptimal scheduler can negate the benefits of a good partition, and vice versa. In this paper, we propose Hyperion, a hierarchical two-stage framework that jointly optimizes partitioning and scheduling to minimize end-to-end latency for pipelined LLM inference in multi-tier networks, balancing compute and memory across tiers while introducing negligible runtime overhead and requiring no model retraining. Motivated by the observation that partition choices evolve on slower timescales than request arrivals, Stage 1 performs offline, inter-tier partitioning via a Hyperion Split with Dynamic Programming (HypSplit-DP) procedure to produce balanced stage times under tier capacity and memory constraints; to adapt to time-varying load, Stage 2 performs online, intra-tier scheduling with a lightweight Hyperion Scheduling for Real-Time (HypSched-RT) algorithm that maps each request to the best available node using real-time estimates of queue length and effective capacity. Experimental results on multi-tier inference tasks demonstrate that Hyperion significantly reduces end-to-end latency by up to 52.1% and 31.2%, with the Phi-3-medium, compared to the GPipe and HEFT baselines, respectively. Furthermore, Hyperion shows superior scalability in long-sequence generation, maintaining a 44.5% lower latency and achieving higher GPU utilization.
@@ I Introduction
To support edge side LLM deployment, we adopt a multi-tier network architecture. As shown in Fig. 1, this integrated multi-tier design comprises devices distributed across three tiers under the central cloud [9]. Through Network AI Management and Orchestration (NAMO), the cloud coordinates cross-domain resources, and orchestrates task placement [10]; tier-level Network AI Logic and Control (NALC) modules execute real-time, region-local optimization. The multi-tier design also offers practical deployability, providing enhanced privacy and flexibility while aligning with geographic or organizational boundaries [11]. For example, in manufacturing, Tier 1 sits at each production line, Tier 2 spans the campus across multiple lines, and Tier 3 resides in the company’s headquarters, connected via secure wireless backhaul.
@@
Therefore, to achieve efficient inference across multi-tier networks, we propose Hyperion, a framework that introduces a two-stage strategy. Crucially, our work moves beyond the isolated, single-tier scheduling common in current research by creating a unified, holistic optimization strategy that considers the entire network hierarchy. To this end, we first formulate a comprehensive Min-Max problem aimed at minimizing the inference latency, denoted as $(\mathcal{P}_{0})$, and then introduce Hierarchical Parallel Execution and resource-aware scheduling for LLM Inference on multi-tier Networks (Hyperion), which decouples the complex joint optimization problem into two less complex, sequential sub-problems. First Stage: Model Partitioning and inter-tier Allocation Strategy $(\mathcal{P}_{1})$., which is an offline static process that partitions the LLM’s blocks across tiers. By considering each tier’s computing and memory resources to optimize load balancing, this stage leverages a Hyperion Split with Dynamic Programming (HypSplit-DP) algorithm to determine the optimal partition strategy. Second Stage: intra-tier Task Scheduling Strategy $(\mathcal{P}_{2})$. Taking the fixed partition from first stage as a given, this stage leverages the Hyperion Scheduling for Real-Time (HypSched-RT) algorithm, which achieves optimal $O(K_{j})$ time complexity via a single linear scan, performing online scheduling of prompts/activations to identify the most suitable intra-layer nodes for pipeline parallel. This hierarchical decoupling strategy significantly reduces the overall computing complexity, providing a scalable and efficient solution. The main contributions of this study are summarized as follows:
@@ III-D Problem Decoupling and Hierarchical Solution
The overall optimization problem, $\boldsymbol{\mathcal{P}_{0}}$, as previously formulated, is a Mixed Integer Nonlinear Programming (MINLP) problem [46]. Its decision variables include integer types (the model partition points $\mathbf{p}$) and binary types (the task scheduling decisions $\mathbf{Y}$), and its objective function contains a nonlinear max operator. Such problems are typically NP hard, and the computing complexity of solving them directly grows dramatically with the model depth $N$ and the number of network nodes, making direct solutions difficult to apply.
@@
  • Sub problem 1: Cross Hierarchy Model Partitioning ($\boldsymbol{\mathcal{P}_{1}}$): This is a inter-tier, static planning problem. Its objective is to find the globally optimal model partition point vector $\mathbf{p}$, under the assumption that each hierarchy can find the optimal node.
@@ IV-A Stage 1: Offline inter-tier LLM Partitioning via HypSplit-DP
Here we aim to determine the optimal model partition vector $\mathbf{p}^{*}=\left(p_{1}^{*},p_{2}^{*},\ldots,p_{T-1}^{*}\right)$, which can be expressed as $\boldsymbol{\mathcal{P}_{1}}$. For notational convenience, we define the computing load of the $j$-th tier as $L_{j}(\mathbf{p})=\frac{1}{C_{j}}\sum_{i=p_{j-1}+1}^{p_{j}}f_{i}$, where $p_{0}=0$ and $p_{T}=N$. Therefore, the objective function can be concisely expressed as $\min_{\mathbf{p}}\max_{j\in\mathcal{T}}\left\{L_{j}(\mathbf{p})\right\}$. It is noteworthy that the volume of data transmitted between adjacent tiers is $\text{batch\_size}\times\text{sequence\_length}\times\text{hidden\_dim}$, which is a constant value independent of the partitioning strategy $\mathbf{p}$. Consequently, the total inter-tier transmission time, represented by $\sum_{j=1}^{T-1}\tau_{j,j+1}$, can be disregarded during the optimization process as it does not affect the determination of the optimal solution.
@@
in which $\vee$ denotes the logical OR operation, $\wedge$ denotes the logical AND operation, and $\cdot$] is the Iverson bracket, which evaluates to 1 if the condition inside is true, and 0 otherwise. To improve computing efficiency, we pre calculate the prefix sums of $f_{i}$ and $m_{i}$:
@@
Hence, the interval sums can be computed efficiently: $\sum_{i=k+1}^{n}f_{i}=S_{f}(n)-S_{f}(k)$ and $\sum_{i=k+1}^{n}m_{i}=S_{m}(n)-S_{m}(k)$. For the base case, at the first tier ($j=1$), the state $DP(1,n)$ is true if and only if placing all of the first $n$ modules on the first tier satisfies the constraints.
@@ V-A Experimental Setup and Baselines
To evaluate the performance of the Hyperion framework, we designed the following network. The setup comprised three computing tiers, each consisting of multiple heterogeneous NVIDIA Jetson nodes. Following the pipeline parallelism model, the inference tasks flow sequentially from Tier 1 to Tier 3. For clarity, the heterogeneity considered in our study is across tiers: within each tier, all nodes are of the same device model. The specific configurations for each tier, including the device type, quantity, computing, and memory, are detailed in Table I. To evaluate the generality of our scheduling framework, we also evaluate Hyperion on the Llama3 8B and Phi-3-medium models, with their architectural differences detailed in Table I. The input size for all tasks was standardized to 64 tokens, and the generation length was set to 128 tokens for each task.
@@ V-B Comparative Analysis of end-to-end Latency
We conducted experiments within the three tier architecture, and the results are shown in Fig. 5. Using the LLaMA3 model, we tested at bandwidths of 1 Gbps, as presented in Fig. 5(a). The x axis represents the number of tasks from 1 to 14, while the y axis shows the average latency required to complete these tasks. The performance of GPipe, HEFT, and Hyperion are represented by the grey dashed line, brown dash dotted line, and blue solid line, respectively. Clearly, the latency of GPipe and HEFT is significantly higher than that of Hyperion, with GPipe consistently showing the highest latency. Specifically, when the task count reaches 14, GPipe and HEFT take approximately 637s and 451s, respectively, while Hyperion takes about 312s. At this point, the latency for GPipe and HEFT are 51.0% and 30.8% higher than Hyperion’s, respectively. This demonstrates that Hyperion effectively implements model partitioning and scheduling, showing its efficiency in the multi-tier network.
In Fig. 5(b), we tested at bandwidths of 100 Mbps. As the number of tasks increased from 1 to 14, the latency for GPipe and HEFT increased from approximately 52.7s to 641.5s and from 38.5s to 471.1s, respectively. This shows total latency increases of 588.8s for GPipe and 432.6s for HEFT. In contrast, Hyperion’s latency increased from about 27.5s to 368.8s, a more modest increase of only 341.3s. This represents a reduction in latency growth of 42.0% compared to GPipe and 21.1% compared to HEFT. This indicates that when the task number surges, Hyperion can optimize task distribution across tiers via its model partitioning algorithm, thereby utilizing resources more efficiently and reducing task latency.
@@
Llama3 8B 1 Gbps Bandwidth Avg. GPU Util. (%) 45.9% 47.2% 52.1%
@@
Table II presents a performance of Hyperion with the Llama3 and Phi-3 models on three-tier network in more detail. Our analysis reveals several key findings. First, the Phi-3-medium model is demonstrably more resource intensive than Llama3 8B, exhibiting higher end-to-end latency and consuming more GPU and memory resources across all configurations. For instance, under the 1 Gbps condition, Phi-3-medium’s latency was 35.9% higher than Llama3’s (38.7s vs. 24.8s), with its memory utilization on the AGX Orin being notably higher (69.2% vs. 42.7%).
@@ V-D Scalability with Increasing Output Tokens
Next, we tested the performance difference between Hyperion and the baseline algorithm under different output tokens. As shown in Fig. 9 (a), GPipe, consistently demonstrates the highest latency, starting at 50.1s for 128 tokens and rising steeply to about 66.2s at 256 tokens. The HEFT algorithm, performs slightly better but follows a similar steep trajectory, with its latency increasing from 35.0s to 48.1s over the same token range. In contrast, Hyperion proves to be the most efficient algorithm. It begins with the lowest latency of approximately 25.2s at 128 tokens and concludes at a significantly lower 35.8s at 256 tokens.
Next we tested Hyperion’s performance on the Phi in Fig. 9 (b), which reveals a widening performance gap as output tokens increase. While the initial latencies at 128 tokens were closely clustered between 40.1s (Hyperion), 53.4s (HEFT) and 78.7s (GPipe), their different scaling efficiencies led to a significant divergence. By the 256 token, Hyperion’s latency reached only 57.6s, whereas HEFT and GPipe climbed to 74.5s and 103.7s, respectively. This demonstrates Hyperion’s superior scalability, making its end-to-end latency approximately 22.7% lower than HEFT’s and 44.5% lower than GPipe’s for longer generative tasks. The consistent outperformance of Hyperion over HEFT and GPipe, suggesting that its resource aware mechanism prevents the creation of “bottleneck” nodes in stage one, thereby optimizing overall latency.
@@ V-E Performance across Different Network Architectures
Our evaluation validates Hyperion’s superiority, demonstrating performance gains over established baselines across multiple models and scenarios. For instance, with the Llama3 8B model under a heavy load of 14 tasks, Hyperion achieved an average latency of approximately 312s, outperforming HEFT by 31.2% and GPipe by 52.1%. The framework’s scalability was particularly evident in long sequence generation tasks. Specifically, when generating 256 output tokens with the Phi-3 model, GPipe’s latency of 103.7s was approximately 44.5% higher than Hyperion’s 57.6s. Moreover, Hyperion’s architecture inherently bolsters data security; by processing data locally at lower tiers, it minimizes network exposure and mitigates interception risks. These findings prove that a hierarchical, resource aware strategy is critical for the practical deployment of LLM at the network edge. This work, therefore, paves the way for more advanced, low latency services, with future research poised to explore additional optimizations such as energy efficiency.

### v2

@@ Abstract
LLMs are increasingly executed in edge where limited GPU memory and heterogeneous computation jointly constrain deployment which motivates model partitioning and request scheduling. In this setting, minimizing latency requires addressing the tight coupling between model placement and request scheduling across heterogeneous nodes, as suboptimal decisions in one domain can negate benefits in the other. In this paper, we propose Hyperion, a hierarchical two-stage framework that jointly optimizes partitioning and scheduling for pipelined LLM inference. Hyperion minimizes latency by balancing resources across tiers without requiring model retraining or incurring significant runtime overhead. Leveraging the timescale difference between partitioning and request arrivals, Stage 1 performs offline, inter-tier partitioning via a Hyperion Split with Dynamic Programming (HypSplit-DP) procedure to produce balanced stage times under tier capacity and memory constraints; to adapt to time-varying load, Stage 2 performs online, intra-tier scheduling with a lightweight Hyperion Scheduling for Real-Time (HypSched-RT) that maps each request to the best available node using real-time estimates of queue length and effective capacity. Experiments with Phi-3-medium demonstrate that Hyperion reduces latency by up to 52.1% (vs. GPipe) and 31.2% (vs. HEFT). Furthermore, Hyperion exhibits superior scalability for long-sequence generation, maintaining 44.5% lower latency and higher GPU utilization.
@@ I Introduction
To facilitate LLM deployment, we adopt a multi-tier network architecture. As shown in Fig. 1, this multi-tier architecture comprises heterogeneous computing nodes distributed across three tiers under the central cloud [9]. Through Network AI Management and Orchestration (NAMO), the cloud coordinates cross-domain resources, and orchestrates task placement [10]; tier-level Network AI Logic and Control (NALC) modules execute real-time, region-local optimization. The multi-tier design also offers practical deployability, providing enhanced privacy and flexibility while aligning with geographic or organizational boundaries [11]. For example, in manufacturing, Tier 1 sits at each production line, Tier 2 spans the campus across multiple lines, and Tier 3 resides in the company’s headquarters, connected via secure wireless backhaul.
@@
Therefore, to achieve efficient inference across multi-tier networks, we propose Hyperion, a framework that introduces a two-stage strategy. Crucially, our work moves beyond the isolated, single-tier scheduling common in current research by creating a unified, holistic optimization strategy that considers the entire network hierarchy. To this end, we first formulate a comprehensive Min-Max problem aimed at minimizing the inference latency, denoted as $(\mathcal{P}_{0})$, and then introduce Hierarchical Parallel Execution and resource-aware scheduling for LLM Inference on multi-tier Networks (Hyperion), which decouples the complex joint optimization problem into two less complex, sequential sub-problems. First Stage: Model Partitioning and inter-tier Allocation Strategy $(\mathcal{P}_{1})$, which is an offline static process that partitions the LLM’s blocks across tiers. By considering each tier’s computing and memory resources to optimize load balancing, this stage leverages a Hyperion Split with Dynamic Programming (HypSplit-DP) algorithm to determine the optimal partition strategy. Second Stage: intra-tier Task Scheduling Strategy $(\mathcal{P}_{2})$. Taking the fixed partition from first stage as a given, this stage leverages the Hyperion Scheduling for Real-Time (HypSched-RT) algorithm, which achieves optimal $O(K_{j})$ time complexity via a single linear scan, performing online scheduling of prompts/activations to identify the most suitable intra-layer nodes for pipeline parallel. This hierarchical decoupling strategy significantly reduces the overall computing complexity, providing a scalable and efficient solution. The main contributions of this study are summarized as follows:
@@ III-D Problem Decoupling and Hierarchical Solution
The overall optimization problem, $\boldsymbol{\mathcal{P}_{0}}$, as previously formulated, can be formulated as a Mixed Integer Nonlinear Programming (MINLP) problem [46]. Its decision variables include integer types (the model partition points $\mathbf{p}$) and binary types (the task scheduling decisions $\mathbf{Y}$), and its objective function contains a nonlinear max operator. Such problems are typically NP hard, and the computing complexity of solving them directly grows dramatically with the model depth $N$ and the number of network nodes, making direct solutions difficult to apply.
@@
  • Sub problem 1: Cross Hierarchy Model Partitioning ($\boldsymbol{\mathcal{P}_{1}}$): This is an inter-tier, static planning problem. Its objective is to find the globally optimal model partition point vector $\mathbf{p}$, under the assumption that each hierarchy can find the optimal node.
@@ IV-A Stage 1: Offline inter-tier LLM Partitioning via HypSplit-DP
We aim to determine the optimal model partition vector $\mathbf{p}^{*}=\left(p_{1}^{*},p_{2}^{*},\ldots,p_{T-1}^{*}\right)$, which can be expressed as $\boldsymbol{\mathcal{P}_{1}}$. For notational convenience, we define the computing load of the $j$-th tier as $L_{j}(\mathbf{p})=\frac{1}{C_{j}}\sum_{i=p_{j-1}+1}^{p_{j}}f_{i}$, where $p_{0}=0$ and $p_{T}=N$. Therefore, the objective function can be concisely expressed as $\min_{\mathbf{p}}\max_{j\in\mathcal{T}}\left\{L_{j}(\mathbf{p})\right\}$. It is noteworthy that the volume of data transmitted between adjacent tiers is $\text{batch\_size}\times\text{sequence\_length}\times\text{hidden\_dim}$, which is a constant value independent of the partitioning strategy $\mathbf{p}$. Consequently, the total inter-tier transmission time, represented by $\sum_{j=1}^{T-1}\tau_{j,j+1}$, can be disregarded during the optimization process as it does not affect the determination of the optimal solution.
@@
in which $\vee$ denotes the logical OR operation, $\wedge$ denotes the logical AND operation, and $\cdot$] is the Iverson bracket, which evaluates to 1 if the condition inside is true, and 0 otherwise. To improve computing efficiency, we precompute the prefix sums of $f_{i}$ and $m_{i}$:
@@
Hence, the interval sums can be computed efficiently: $\sum_{i=k+1}^{n}f_{i}=S_{f}(n)-S_{f}(k)$ and $\sum_{i=k+1}^{n}m_{i}=S_{m}(n)-S_{m}(k)$. For the base case, at the first tier ($j=1$), the state $DP(1,n)$ is true if and only if placing all of the first $n$ modules on the first tier satisfies the constraints:
@@ V-A Experimental Setup and Baselines
To evaluate the performance of the Hyperion framework, we designed the following network. The setup comprised three computing tiers, each consisting of multiple heterogeneous NVIDIA Jetson nodes. Following the pipeline parallelism model, the inference tasks flow sequentially from Tier 1 to Tier 3. For clarity, the heterogeneity considered in our study is across tiers: within each tier, all nodes are of the same device model. The specific configurations for each tier, including the device type, quantity, computing, and memory, are detailed in Table I. To evaluate the generality of our scheduling framework, we also evaluate Hyperion on the Llama3-8B and Phi-3-medium models, with their architectural differences detailed in Table I. The input size for all tasks was standardized to 64 tokens, and the generation length was set to 128 tokens for each task.
@@ V-B Comparative Analysis of end-to-end Latency
We conduct experiments within the three tier architecture, and the results are shown in Fig. 5. Using the LLaMA3 model, we test at bandwidths of 1 Gbps, as presented in Fig. 5(a). The x axis represents the number of tasks from 1 to 14, while the y axis shows the average latency required to complete these tasks. The performance of GPipe, HEFT, and Hyperion are represented by the grey dashed line, brown dash dotted line, and blue solid line, respectively. Clearly, the latency of GPipe and HEFT is significantly higher than that of Hyperion, with GPipe consistently showing the highest latency. Specifically, when the task count reaches 14, GPipe and HEFT take approximately 637s and 451s, respectively, while Hyperion takes about 312s. At this point, the latency for GPipe and HEFT are 51.0% and 30.8% higher than Hyperion’s, respectively. This demonstrates that Hyperion effectively implements model partitioning and scheduling, showing its efficiency in the multi-tier network.
In Fig. 5(b), we test at bandwidths of 100 Mbps. As the number of tasks increased from 1 to 14, the latency for GPipe and HEFT increased from approximately 52.7s to 641.5s and from 38.5s to 471.1s, respectively. This shows total latency increases of 588.8s for GPipe and 432.6s for HEFT. In contrast, Hyperion’s latency increased from about 27.5s to 368.8s, a more modest increase of only 341.3s. This represents a reduction in latency growth of 42.0% compared to GPipe and 21.1% compared to HEFT. This indicates that when the task number surges, Hyperion can optimize task distribution across tiers via its model partitioning algorithm, thereby utilizing resources more efficiently and reducing task latency.
@@
Llama3-8B 1 Gbps Bandwidth Avg. GPU Util. (%) 45.9% 47.2% 52.1%
@@
Table II presents a performance of Hyperion with the Llama3 and Phi-3 models on three-tier network in more detail. Our analysis reveals several key findings. First, the Phi-3-medium model is demonstrably more resource intensive than Llama3-8B, exhibiting higher end-to-end latency and consuming more GPU and memory resources across all configurations. For instance, under the 1 Gbps condition, Phi-3-medium’s latency was 35.9% higher than Llama3’s (38.7s vs. 24.8s), with its memory utilization on the AGX Orin being notably higher (69.2% vs. 42.7%).
@@ V-D Scalability with Increasing Output Tokens
Next, we test the performance difference between Hyperion and the baseline algorithm under different output tokens. As shown in Fig. 9 (a), GPipe, consistently demonstrates the highest latency, starting at 50.1s for 128 tokens and rising steeply to about 66.2s at 256 tokens. The HEFT algorithm, performs slightly better but follows a similar steep trajectory, with its latency increasing from 35.0s to 48.1s over the same token range. In contrast, Hyperion proves to be the most efficient algorithm. It begins with the lowest latency of approximately 25.2s at 128 tokens and concludes at a significantly lower 35.8s at 256 tokens.
Next we test Hyperion’s performance on the Phi in Fig. 9 (b), which reveals a widening performance gap as output tokens increase. While the initial latencies at 128 tokens were closely clustered between 40.1s (Hyperion), 53.4s (HEFT) and 78.7s (GPipe), their different scaling efficiencies led to a significant divergence. By the 256 token, Hyperion’s latency reached only 57.6s, whereas HEFT and GPipe climbed to 74.5s and 103.7s, respectively. This demonstrates Hyperion’s superior scalability, making its end-to-end latency approximately 22.7% lower than HEFT’s and 44.5% lower than GPipe’s for longer generative tasks. The consistent outperformance of Hyperion over HEFT and GPipe, suggesting that its resource aware mechanism prevents the creation of “bottleneck” nodes in stage one, thereby optimizing overall latency.
@@ V-E Performance across Different Network Architectures
Our evaluation validates Hyperion’s superiority, demonstrating performance gains over established baselines across multiple models and scenarios. For instance, with the Llama3-8B model under a heavy load of 14 tasks, Hyperion achieved an average latency of approximately 312s, outperforming HEFT by 31.2% and GPipe by 52.1%. The framework’s scalability was particularly evident in long sequence generation tasks. Specifically, when generating 256 output tokens with the Phi-3 model, GPipe’s latency of 103.7s was approximately 44.5% higher than Hyperion’s 57.6s. Moreover, Hyperion’s architecture inherently bolsters data security; by processing data locally at lower tiers, it minimizes network exposure and mitigates interception risks. These findings prove that a hierarchical, resource aware strategy is critical for the practical deployment of LLM at the network edge. This work, therefore, paves the way for more advanced, low latency services, with future research poised to explore additional optimizations such as energy efficiency.